OpenStudy (anonymous):

A 12-foot ladder is placed 4 feet from the wall. Approximately how far up the wall will the ladder reach? 11 16 8

OpenStudy (nikato):

You can use the pythagoreom theorem

OpenStudy (anonymous):

11

OpenStudy (nikato):



OpenStudy (nikato):

x^2 + 4^2=12^2

OpenStudy (anonymous):

is it 11 @nikato

OpenStudy (nikato):

yes. x^2 + 16=144 -16 -16 ------------ x^2 = 128 x is about 11

OpenStudy (anonymous):

but cant it be 8 @nikato

OpenStudy (anonymous):

cuz u do 128 divide 16=8

OpenStudy (nikato):

Why would you divide 128 by 16?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

but how did u get 11??

OpenStudy (nikato):

So you have the pythagoreom theorem based on that diagram I drew. x^2 + 4^2 =12^2 x^2 +16 =144. Subtract 16 on both sides x^2=128. Now square root both sides x= sqrt 128 x= 11.313.... So x=11. By rounding

OpenStudy (nikato):

Does it make sense now? Oh I plugged in sqrt 128 in the calculator to get that number in case you're wondering

OpenStudy (anonymous):

oh I get it thx :)

OpenStudy (nikato):

No problem

OpenStudy (anonymous):

but wait how did u get x= 11.313...

OpenStudy (nikato):

Calculator... Square root of 128

OpenStudy (nikato):

so if you have x^2, in order to get x by itself,